2. Consider the reaction

4PH3(g) → P4(g) + 6H2(g)

Suppose that, at a particular moment during the reaction, molecular hydrogen is being formed at rate of 0.234 M/s.

(a) At what rate is P4 changing?
_______ M/s

(b) At what rate is PH3 changing?
________ M/sF
Answer #1

4-PH3(g) → P4 g) + 6H2(g) A1H2) = 0.234 m/s st rate of reeaction = 1 7 A[PH₂) = A[P4] J-1 H2) ot (0) Or P4) set = _ACHE 4 x 0